# Multigranular Evaluation for Brain Visual Decoding

### v1

@@ Appendix A Details on Stimulus-Neuroimaging Datasets
This section provides details on stimulus-neuroimaging datasets. Tab. S1 summarizes the core properties of these stimulus-neuroimaging combinations. The Natural Scenes Dataset allen2022massive has recently garnered significant attention, with many methods proposed for its analysis, and most of our experiments focus on methods applied to this dataset for comprehensive and detailed analysis.
@@
NSD allen2022massive is the largest public 7 Tesla fMRI scans, featuring brain recordings from eight participants who passively viewed images from the Common Objects in Context (COCO) dataset lin2014microsoft for up to 40 hours in an MRI machine. Each image was shown for three seconds and repeated three times across 30-40 scanning sessions, resulting in 22,000-30,000 fMRI response trials per participant.
@@
EEG-Things grootswagers2022human provides neuroimaging recordings for a systematic collection of objects and concepts, featuring electroencephalography (EEG) responses from 50 subjects to a stimulus set of 1,854 object concepts and 22,248 images.
@@
CC2017 wen2018neural is a public benchmark consisting of video clips and corresponding fMRI recordings collected from three subjects using a 3T MRI scanner. The training set includes 18 segments of 8-minute video clips (totaling 2.4 hours), yielding 4,320 paired fMRI-video examples. The test set includes 5 segments (totaling 40 minutes), providing 1,200 samples.
@@
SEED-DV liu2024eeg2video includes EEG recordings from 20 subjects using a 62-channel system as they viewed 1,400 video clips covering 40 object concepts. These concepts were grouped into 9 broader categories, offering dynamic visual stimuli paired with corresponding neural responses.
@@
The fMRI-3D dataset comprises two successive components for neural decoding of 3D visual perception. fMRI-Shape gao2024mind3d collects fMRI responses from 14 participants viewing 360-degree videos of 1,624 3D objects spanning 55 categories, totaling 123,200 frames. The subsequent fMRI-Objaverse gao2024mind3d++ further expands category diversity and object coverage, with data from 5 participants exposed to 3,142 3D objects spanning 117 categories, each accompanied by descriptive text captions.
@@
EEG-3D guo2025neuro3d provides extensive EEG recordings from 12 participants who viewed 3D objects across 72 categories, rendered as both videos and images. Each participant underwent approximately 5.5 hours of recording. The dataset includes 10 objects per category, selected for shape diversity, with each object paired with a descriptive text caption.
@@
NSD allen2022massive Nature Neuroscience 2022 fMRI Image takagi2023improving ; ozcelik2023brain ; scotti2023reconstructing ; scotti2024mindeye2 ; xia2024dream ; xia2024umbrae ; shen2024neuro ; huo2024neuropictor ; gong2025mindtuner ; liu2025see
EEG-Things grootswagers2022human Scientific Data 2022 EEG Image li2024visual ; zhang2025cognitioncapturer
CC2017 wen2018neural Cerebral Cortex 2018 fMRI Video chen2023cinematic ; li2025decofuse ; gong2024neuroclips
SEED-DV liu2024eeg2video NeurIPS 2024 EEG Video liu2024eeg2video
fMRI-3D gao2024mind3d ECCV 2024 fMRI 3D Shape gao2024mind3d ; gao2024mind3d++
EEG-3D guo2025neuro3d CVPR 2025 EEG 3D Shape guo2025neuro3d
The latter three datasets, SEED-DV liu2024eeg2video , fMRI-3D gao2024mind3d , EEG-3D guo2025neuro3d , are the most recently proposed datasets, with limited comparisons available. We share their results to demonstrate that the proposed evaluation framework can be extended to different stimulus-neuroimaging datasets.
@@ Appendix B Details on Prior Evaluation Procedure
Tab. S2 presents results on NSD and EEG-Things using eight commonly used metrics: PixCorr, SSIM wang2004image , AlexNet krizhevsky2017imagenet -2/5, Inception szegedy2016rethinking , CLIP radford2021learning , EffNet tan2019efficientnet , and SwAV caron2020unsupervised . PixCorr measures the Pearson correlation between the pixel values of the ground-truth (GT) image and the reconstruction, quantifying the pixel-level similarity. SSIM (Structural Similarity Index) wang2004image compares the structural and textural features of the GT and the reconstruction. The AlexNet-2/5 metrics refer to the similarity between the embeddings of the GT and the reconstruction, based on the deep features extracted from the second and fifth layers of the pretrained AlexNet krizhevsky2017imagenet , respectively. Similarly, Inception refers to the comparison of the final pooling layer in InceptionV3 szegedy2016rethinking , and CLIP corresponds to the comparison of embeddings from the final layer of CLIP-Vision radford2021learning . EffNet-B and SwAV metrics refer to the distance or similarity between the GT embedding and the reconstructed embedding using EfficientNet-B1 tan2019efficientnet and SwAV-ResNet50 caron2020unsupervised , respectively.
Tab. S3 presents a quantitative evaluation of video reconstruction on CC2017 wen2018neural , using 2-way and 50-way semantic accuracy as the semantic metrics, and the ratio of foreground-background matching (mask matching ratio, MMR) between the ground truth and decoded images as the spatial-level metric. Tab. S5, Tab. S5, and Tab. S6 shows quantitative evaluation of visual reconstruction on SEED-DV liu2024eeg2video , EEG-3D guo2025neuro3d , and fMRI-3D gao2024mind3d , with results excerpted from the corresponding papers. For ease of comparison, we also include BASIC-L and BASIC-H evaluation results. As shown, while NSD metrics saturates in scores and fails to distinguish method performance, our metrics, BASIC-L and BASIC-H, provide clearer differentiation and more discriminative information across objects, attributes, and relations for high-level semantics, as well as salient, semantic, instance, and part granularity for low-level structure. Details for each dimension are in Tab. 2 and Tab. 3.
@@
Other modality-specific metrics are also introduced. For instance, fMRI-3D evaluation uses common 3D reconstruction metrics, including Fréchet Point Cloud Distance (FPD) (scaled by ×$\times 10^{-1}$), Chamfer Distance (CD) (scaled by ×$\times 10^{2}$), and Earth Mover’s Distance (EMD) (scaled by ×$\times 10^{2}$). These metrics are computed by sampling point clouds from both the ground truth and the generated meshes.
Table S2: Quantitative evaluation for fMRI-to-image reconstruction on NSD allen2022massive and EEG-to-image reconstruction on EEG-Things grootswagers2022human following standard NSD and our BASIC metrics. While NSD metrics saturates in scores and fails to distinguish method performance, our metrics provide clearer differentiation and more discriminative information. See detailed sub-scores in Tab. 2 and Tab. 3.
@@
NSD  allen2022massive
SDRecon takagi2023improving 0.246 0.410 78.9% 85.6% 11.81 83.8% 82.1% 0.811 0.504 35.31
BrainDiffuser ozcelik2023brain 0.273 0.365 94.4% 96.6% 16.65 91.3% 90.9% 0.728 0.422 39.71
MindEye scotti2023reconstructing 0.319 0.360 92.8% 96.9% 17.03 94.6% 93.3% 0.648 0.377 44.30
DREAM xia2024dream 0.274 0.328 93.9% 96.7% 19.57 93.4% 94.1% 0.645 0.418 46.37
MindEye2 scotti2024mindeye2 0.322 0.431 96.1% 98.6% 22.16 95.4% 93.0% 0.619 0.344 44.39
MindBridge wang2024mindbridge 0.151 0.263 87.7% 95.5% 15.00 92.4% 94.7% 0.712 0.418 40.16
UMBRAE xia2024umbrae 0.283 0.341 95.5% 97.0% 17.89 91.7% 93.5% 0.700 0.393 44.06
NeuroPictor huo2024neuropictor 0.229 0.375 96.5% 98.4% 25.88 94.5% 93.3% 0.639 0.350 44.21
NeuroVLA shen2024neuro 0.265 0.357 93.1% 97.1% 13.54 96.8% 97.5% 0.633 0.321 47.88
SepBrain wang2024unibrain 0.309 0.317 94.2% 97.4% 18.84 94.5% 95.3% 0.656 0.374 43.04
UniBrain wang2024unibrain 0.155 0.259 87.8% 95.9% 13.79 92.4% 94.0% 0.691 0.407 39.89
STTM liu2025see 0.333 0.334 95.7% 98.5% 22.90 95.8% 95.7% 0.611 0.338 45.88
MindTuner gong2025mindtuner 0.322 0.421 95.8% 98.8% 16.98 95.6% 93.8% 0.612 0.340 44.63
BrainGuard tian2025brainguard 0.313 0.330 94.7% 97.8% 21.76 96.1% 96.4% 0.624 0.353 45.43
EEG-Things  grootswagers2022human
ATM li2024visual 0.160 0.345 0.776 0.866 17.60 0.734 0.786 - 0.582 30.55
CogCapturer zhang2025cognitioncapturer 0.175 0.366 0.760 0.610 16.22 0.721 0.744 - 0.577 28.64
Table S3: Quantitative evaulation for fMRI-to-video reconstruction on CC2017 wen2018neural .
@@
MinD-Video chen2023cinematic 0.792 0.172 28.63 0.660 15.25
NeuroClips gong2024neuroclips 0.808 0.195 45.12 0.687 23.52
DecoFuse li2025decofuse 0.824 0.208 29.03 0.706 15.31
Table S4: Quantitative evaluation for EEG-to-video reconstruction on SEED-DV liu2024eeg2video .
@@
EEG2Video liu2024eeg2video 0.798 0.159 49.54 0.774 0.138 0.256 20.54
Table S5: Quantitative evaluation for EEG-to-3D reconstruction on EEG-3D guo2025neuro3d .
@@
Neuro-3D guo2025neuro3d 55.81 35.89 72.08 57.64 9.69 23.08
Table S6: Quantitative evaluation for fMRI-to-3D reconstruction on fMRI-3D gao2024mind3d ; gao2024mind3d++ .
@@
MinD-3D gao2024mind3d fMRI-Shape 0.828 0.459 3.157 1.742 3.833 0.306 32.81 0.674 14.72 30.95
MinD-3D++ gao2024mind3d++ 0.887 0.616 3.025 1.635 3.672 0.234 34.09 0.763 12.62 35.08
MinD-3D gao2024mind3d fMRI-Objaverse 0.793 0.427 4.304 2.142 5.323 0.544 31.09 0.724 N/A N/A
MinD-3D++ gao2024mind3d++ 0.894 0.618 3.325 1.779 4.073 0.343 33.64 0.808 N/A N/A
@@ Appendix C Details on BASIC
This section complements Sec. 3 with additional implementation details and qualitative results.
C.1 Structured Entity Extraction from Captions
@@
We adopt LLaVA-1.6-13B liu2023visual as our default MLLM, considering the trade-off between computational efficiency and captioning accuracy. Recent studies on detailed captioning evaluation lu2024benchmarking ; dong2024benchmarking suggest that this configuration achieves top-tier performance in long image description, comparable to human judgments. For prompting, we adopt a simple instruction— “Describe the image in detail”—instead of crafting elaborate prompts that explicitly reflect evaluation dimensions outlined in Tab. 1. Preliminary experiments indicate that this simple prompt is sufficient to elicit detailed captions covering the desired evaluation dimensions when using LLaVA liu2023visual . This may be attributed to LLaVA’s training corpus, which likely includes similarly phrased instructions for detailed image captioning. Further discussions on the qualitative and quantitative effects of different prompts and MLLMs are presented in the following sections.
@@
    "A peaceful beach with soft white sand stretching along the coastline, where turquoise ocean waves gently roll onto the shore. Several people are sunbathing near the water while others are playing volleyball in the distance."
@@
    "A tropical island beach with lush palm trees swaying in the breeze, and the bright blue sea sparkling under the sun. Some people are lounging under umbrellas by the water, while a group of friends is playing volleyball near the edge of the beach."
@@
Figure S1: The visualization of multigranular segmentation. Reference images and structured annotations can be found in Fig. S2 and Tab. S8, respectively.
C.2 Multigranular Image Segmentation
This section provides additional implementation details and qualitative results to complement the discussion in Sec. 3.3. For clarity and consistency, we follow the same two subtitles.
@@ C.2 Multigranular Image Segmentation
Multigranular segmentation involves decomposing visual scenes into hierarchical levels of detail, from whole-object recognition to fine-grained part segmentation. We first identify multigranular semantic instances that serve as textual instructions for subsequent segmentation using a multimodal large language model liu2023visual . The full categorization prompt used for multigranular object identification is provided in Tab. S7. As illustrated in Tab. S8, each image is annotated with structured information, including foreground objects, their constituent parts, and relevant contextual background elements. This hierarchical annotation schema allows for fine-to-coarse segmentation and supports comprehensive evaluation of vision models in terms of object-level recognition and detailed part-level understanding across diverse scene contexts.
@@
Using the structured annotations described above as natural language prompts, we apply Grounded-SAM2 ren2024grounded to produce multigranular segmentation masks. Prompts such as “zebra”, “tail”, or “legs” lead the model to localize and segment corresponding entities and component within an image. This enables model evaluation on fine-grained visual grounding, compositional understanding, and spatial reasoning. The resulting multigranular segmentation visualizations are shown in Fig. S1, with corresponding reference images displayed in Fig. S2 and structured annotations provided in Tab. S8. Taking the top-right three images of the airplane as an example, we present the salient foreground (airplane), background (sky, runway, grass, and trees), and the components of the airplane (fuselage, wings, engines, and tail), respectively.
@@
You are a visual understanding assistant. Given an input image, please analyze and describe it with a structured categorization. Each response should include:
1. Foreground Objects: List the main objects in the foreground and their semantic categories.
2. Background Elements: Describe elements in the background and their semantic categories.
3. Part-level Categories: For each foreground object, identify the visible parts in a general way (e.g., "car" $\rightarrow$ "wheels", "doors").
Provide your answer in a structured JSON format like this:
@@
Table S8: Structured scene annotations for four image instances. ID 0 to 3 corresponds to references in Fig. S2. Each scene is decomposed into foreground (F) objects (e.g., boats, airplane, zebra), their associated part-level components, and background (B) elements (e.g., trees, sky, runway) that represent the broader environmental context. This is derived from a standardized JSON format.

### v2

@@ Appendix A Details on Stimulus-Neuroimaging Datasets
This section provides details on stimulus-neuroimaging datasets. Table S1 summarizes the core properties of these stimulus-neuroimaging combinations. The Natural Scenes Dataset (Allen et al. 2022) has recently garnered significant attention, with many methods proposed for its analysis, and most of our experiments focus on methods applied to this dataset for comprehensive and detailed analysis.
@@
NSD (Allen et al. 2022) is the largest public 7 Tesla fMRI scans, featuring brain recordings from eight participants who passively viewed images from the Common Objects in Context (COCO) dataset (Lin et al. 2014) for up to 40 hours in an MRI machine. Each image was shown for three seconds and repeated three times across 30-40 scanning sessions, resulting in 22,000-30,000 fMRI response trials per participant.
@@
EEG-Things (Grootswagers et al. 2022) provides neuroimaging recordings for a systematic collection of objects and concepts, featuring electroencephalography (EEG) responses from 50 subjects to a stimulus set of 1,854 object concepts and 22,248 images.
@@
CC2017 (Wen et al. 2018) is a public benchmark consisting of video clips and corresponding fMRI recordings collected from three subjects using a 3T MRI scanner. The training set includes 18 segments of 8-minute video clips (totaling 2.4 hours), yielding 4,320 paired fMRI-video examples. The test set includes 5 segments (totaling 40 minutes), providing 1,200 samples.
@@
SEED-DV (Liu et al. 2024b) includes EEG recordings from 20 subjects using a 62-channel system as they viewed 1,400 video clips covering 40 object concepts. These concepts were grouped into 9 broader categories, offering dynamic visual stimuli paired with corresponding neural responses.
@@
The fMRI-3D dataset comprises two successive components for neural decoding of 3D visual perception. fMRI-Shape (Gao et al. 2024) collects fMRI responses from 14 participants viewing 360-degree videos of 1,624 3D objects spanning 55 categories, totaling 123,200 frames. The subsequent fMRI-Objaverse (Gao et al. 2025) further expands category diversity and object coverage, with data from 5 participants exposed to 3,142 3D objects spanning 117 categories, each accompanied by descriptive text captions.
@@
EEG-3D (Guo et al. 2025) provides extensive EEG recordings from 12 participants who viewed 3D objects across 72 categories, rendered as both videos and images. Each participant underwent approximately 5.5 hours of recording. The dataset includes 10 objects per category, selected for shape diversity, with each object paired with a descriptive text caption.
@@
NSD (Allen et al. 2022) Nature Neuroscience 2022 fMRI Image Takagi and Nishimoto (2023b); Scotti et al. (2023, 2024)
EEG-Things (Grootswagers et al. 2022) Scientific Data 2022 EEG Image Li et al. (2024); Zhang et al. (2025)
CC2017 (Wen et al. 2018) Cerebral Cortex 2018 fMRI Video Li et al. (2025); Gong et al. (2024)
SEED-DV (Liu et al. 2024b) NeurIPS 2024 EEG Video Liu et al. (2024b)
fMRI-3D (Gao et al. 2024) ECCV 2024 fMRI 3D Shape Gao et al. (2024, 2025)
EEG-3D (Guo et al. 2025) CVPR 2025 EEG 3D Shape Guo et al. (2025)
The latter three datasets, SEED-DV (Liu et al. 2024b), fMRI-3D (Gao et al. 2024), EEG-3D (Guo et al. 2025), are the most recently proposed datasets, with limited comparisons available. We share their results to demonstrate that the proposed evaluation framework can be extended to different stimulus-neuroimaging datasets.
@@ Appendix B Details on Prior Evaluation Procedure
Table S2 presents results on NSD and EEG-Things using eight commonly used metrics: PixCorr, SSIM (Wang et al. 2004), AlexNet (Krizhevsky, Sutskever, and Hinton 2017)-2/5, Inception (Szegedy et al. 2016), CLIP (Radford et al. 2021), EffNet (Tan and Le 2019), and SwAV (Caron et al. 2020). PixCorr measures the Pearson correlation between the pixel values of the ground-truth (GT) image and the reconstruction, quantifying the pixel-level similarity. SSIM (Structural Similarity Index) (Wang et al. 2004) compares the structural and textural features of the GT and the reconstruction. The AlexNet-2/5 metrics refer to the similarity between the embeddings of the GT and the reconstruction, based on the deep features extracted from the second and fifth layers of the pretrained AlexNet (Krizhevsky, Sutskever, and Hinton 2017), respectively. Similarly, Inception refers to the comparison of the final pooling layer in InceptionV3 (Szegedy et al. 2016), and CLIP corresponds to the comparison of embeddings from the final layer of CLIP-Vision (Radford et al. 2021). EffNet-B and SwAV metrics refer to the distance or similarity between the GT embedding and the reconstructed embedding using EfficientNet-B1 (Tan and Le 2019) and SwAV-ResNet50 (Caron et al. 2020), respectively.
Table S3 presents a quantitative evaluation of video reconstruction on CC2017 (Wen et al. 2018), using 2-way and 50-way semantic accuracy as the semantic metrics, and the ratio of foreground-background matching (mask matching ratio, MMR) between the ground truth and decoded images as the spatial-level metric. Table S5, Table S5, and Table S6 shows quantitative evaluation of visual reconstruction on SEED-DV (Liu et al. 2024b), EEG-3D (Guo et al. 2025), and fMRI-3D (Gao et al. 2024), with results excerpted from the corresponding papers. For ease of comparison, we also include BASIC-L and BASIC-H evaluation results. As shown, while NSD metrics saturates in scores and fails to distinguish method performance, our metrics, BASIC-L and BASIC-H, provide clearer differentiation and more discriminative information across objects, attributes, and relations for high-level semantics, as well as salient, semantic, instance, and part granularity for low-level structure. Details for each dimension are in Table 2 and Table 3.
@@
Other modality-specific metrics are also introduced. For instance, fMRI-3D evaluation uses common 3D reconstruction metrics, including Fréchet Point Cloud Distance (FPD) (scaled by $\times 10^{-1}$), Chamfer Distance (CD) (scaled by $\times 10^{2}$), and Earth Mover’s Distance (EMD) (scaled by $\times 10^{2}$). These metrics are computed by sampling point clouds from both the ground truth and the generated meshes.
Table S2: Quantitative evaluation for fMRI-to-image reconstruction on NSD (Allen et al. 2022) and EEG-to-image reconstruction on EEG-Things (Grootswagers et al. 2022) following standard NSD and our BASIC metrics. While NSD metrics saturates in scores and fails to distinguish method performance, our metrics provide clearer differentiation and more discriminative information. See detailed sub-scores in Table 2 and Table 3.
@@
NSD  (Allen et al. 2022)
SDRecon (Takagi and Nishimoto 2023b) 0.246 0.410 78.9% 85.6% 11.81 83.8% 82.1% 0.811 0.504 35.31
BrainDiffuser (Ozcelik and VanRullen 2023) 0.273 0.365 94.4% 96.6% 16.65 91.3% 90.9% 0.728 0.422 39.71
MindEye (Scotti et al. 2023) 0.319 0.360 92.8% 96.9% 17.03 94.6% 93.3% 0.648 0.377 44.30
DREAM (Xia et al. 2024b) 0.274 0.328 93.9% 96.7% 19.57 93.4% 94.1% 0.645 0.418 46.37
MindEye2 (Scotti et al. 2024) 0.322 0.431 96.1% 98.6% 22.16 95.4% 93.0% 0.619 0.344 44.39
MindBridge (Wang et al. 2024a) 0.151 0.263 87.7% 95.5% 15.00 92.4% 94.7% 0.712 0.418 40.16
UMBRAE (Xia et al. 2024a) 0.283 0.341 95.5% 97.0% 17.89 91.7% 93.5% 0.700 0.393 44.06
NeuroPictor (Huo et al. 2024) 0.229 0.375 96.5% 98.4% 25.88 94.5% 93.3% 0.639 0.350 44.21
NeuroVLA (Shen et al. 2024) 0.265 0.357 93.1% 97.1% 13.54 96.8% 97.5% 0.633 0.321 47.88
SepBrain (Wang et al. 2024b) 0.309 0.317 94.2% 97.4% 18.84 94.5% 95.3% 0.656 0.374 43.04
UniBrain (Wang et al. 2024b) 0.155 0.259 87.8% 95.9% 13.79 92.4% 94.0% 0.691 0.407 39.89
STTM (Liu et al. 2025) 0.333 0.334 95.7% 98.5% 22.90 95.8% 95.7% 0.611 0.338 45.88
MindTuner (Gong et al. 2025) 0.322 0.421 95.8% 98.8% 16.98 95.6% 93.8% 0.612 0.340 44.63
BrainGuard (Tian et al. 2025) 0.313 0.330 94.7% 97.8% 21.76 96.1% 96.4% 0.624 0.353 45.43
EEG-Things  (Grootswagers et al. 2022)
ATM (Li et al. 2024) 0.160 0.345 0.776 0.866 17.60 0.734 0.786 - 0.582 30.55
CogCapturer (Zhang et al. 2025) 0.175 0.366 0.760 0.610 16.22 0.721 0.744 - 0.577 28.64
Table S3: Quantitative evaulation for fMRI-to-video reconstruction on CC2017 (Wen et al. 2018).
@@
MinD-Video (Chen, Qing, and Zhou 2023) 0.792 0.172 28.63 0.660 15.25
NeuroClips (Gong et al. 2024) 0.808 0.195 45.12 0.687 23.52
DecoFuse (Li et al. 2025) 0.824 0.208 29.03 0.706 15.31
Table S4: Quantitative evaluation for EEG-to-video reconstruction on SEED-DV (Liu et al. 2024b).
@@
EEG2Video (Liu et al. 2024b) 0.798 0.159 49.54 0.774 0.138 0.256 20.54
Table S5: Quantitative evaluation for EEG-to-3D reconstruction on EEG-3D (Guo et al. 2025).
@@
Neuro-3D (Guo et al. 2025) 55.81 35.89 72.08 57.64 9.69 23.08
Table S6: Quantitative evaluation for fMRI-to-3D reconstruction on fMRI-3D (Gao et al. 2024, 2025).
@@
MinD-3D (Gao et al. 2024) fMRI-Shape 0.828 0.459 3.157 1.742 3.833 0.306 32.81 0.674 14.72 30.95
MinD-3D++ (Gao et al. 2025) 0.887 0.616 3.025 1.635 3.672 0.234 34.09 0.763 12.62 35.08
MinD-3D (Gao et al. 2024) fMRI-Objaverse 0.793 0.427 4.304 2.142 5.323 0.544 31.09 0.724 N/A N/A
MinD-3D++ (Gao et al. 2025) 0.894 0.618 3.325 1.779 4.073 0.343 33.64 0.808 N/A N/A
@@ Appendix C Details on BASIC
This section complements the method section in the main paer with additional implementation details and qualitative results.
Structured Entity Extraction from Captions
@@
We adopt LLaVA-1.6-13B (Liu et al. 2023) as our default MLLM, considering the trade-off between computational efficiency and captioning accuracy. Recent studies on detailed captioning evaluation (Lu et al. 2025; Dong et al. 2024) suggest that this configuration achieves top-tier performance in long image description, comparable to human judgments. For prompting, we adopt a simple instruction— “Describe the image in detail”—instead of crafting elaborate prompts that explicitly reflect evaluation dimensions outlined in Table 1. Preliminary experiments indicate that this simple prompt is sufficient to elicit detailed captions covering the desired evaluation dimensions when using LLaVA (Liu et al. 2023) . This may be attributed to LLaVA’s training corpus, which likely includes similarly phrased instructions for detailed image captioning. Further discussions on the qualitative and quantitative effects of different prompts and MLLMs are presented in the following sections.
@@
    ”A peaceful beach with soft white sand stretching along the coastline, where turquoise ocean waves gently roll onto the shore. Several people are sunbathing near the water while others are playing volleyball in the distance.”
@@
    ”A tropical island beach with lush palm trees swaying in the breeze, and the bright blue sea sparkling under the sun. Some people are lounging under umbrellas by the water, while a group of friends is playing volleyball near the edge of the beach.”
@@
Figure S1: The visualization of multigranular segmentation. Reference images and structured annotations can be found in Fig. S2 and Table S8, respectively.
Multigranular Image Segmentation
This section provides additional implementation details and qualitative results to complement the main paper.
@@
Multigranular segmentation involves decomposing visual scenes into hierarchical levels of detail, from whole-object recognition to fine-grained part segmentation. We first identify multigranular semantic instances that serve as textual instructions for subsequent segmentation using a multimodal large language model (Liu et al. 2023). The full categorization prompt used for multigranular object identification is provided in Table S7. As illustrated in Table S8, each image is annotated with structured information, including foreground objects, their constituent parts, and relevant contextual background elements. This hierarchical annotation schema allows for fine-to-coarse segmentation and supports comprehensive evaluation of vision models in terms of object-level recognition and detailed part-level understanding across diverse scene contexts.
@@
Using the structured annotations described above as natural language prompts, we apply Grounded-SAM2 (Ren et al. 2023) to produce multigranular segmentation masks. Prompts such as “zebra”, “tail”, or “legs” lead the model to localize and segment corresponding entities and component within an image. This enables model evaluation on fine-grained visual grounding, compositional understanding, and spatial reasoning. The resulting multigranular segmentation visualizations are shown in Fig. S1, with corresponding reference images displayed in Fig. S2 and structured annotations provided in Table S8. Taking the top-right three images of the airplane as an example, we present the salient foreground (airplane), background (sky, runway, grass, and trees), and the components of the airplane (fuselage, wings, engines, and tail), respectively.
@@
You are a visual understanding assistant. Given an input image, please analyze and describe it with a structured categorization. Each response should include: 1. Foreground Objects: List the main objects in the foreground and their semantic categories. 2. Background Elements: Describe elements in the background and their semantic categories. 3. Part-level Categories: For each foreground object, identify the visible parts in a general way (e.g., ”car” $\rightarrow$ ”wheels”, ”doors”). Provide your answer in a structured JSON format like this:
@@
Table S8: Structured scene annotations for four image instances. ID 0 to 3 corresponds to references in Fig. S2. Each scene is decomposed into foreground (F) objects (e.g., boats, airplane, zebra), their associated part-level components, and background (B) elements (e.g., trees, sky, runway) that represent the environmental context. This is from a standardized JSON format.